# Stability and convergence of multi-product expansion splitting methods with negative weights for semilinear parabolic equations

### v1

@@ Abstract
The operator splitting method has been widely used to solve differential equations by splitting the equation into more manageable parts. In this work, we resolves a long-standing problem—how to establish the stability of multi-product expansion (MPE) splitting methods with negative weights. The difficulty occurs because negative weights in high-order MPE method cause the sum of the absolute values of weights larger than one, making standard stability proofs fail. In particular, we take the semilinear parabolic equation as a typical model and establish the stability of arbitrarily high-order MPE splitting methods with positive time steps but possibly negative weights. Rigorous convergence analysis is subsequently obtained from the stability result. Extensive numerical experiments validate the stability and accuracy of various high-order MPE splitting methods, highlighting their efficiency and robustness.
@@ 1 Introduction
Operator splitting method plays a critical role in approximating solutions of differential equations involving the coupling of multiple physical fields. The core idea is to subdivide the origin equations into a series of simpler problems, each corresponding to a different physical or mathematical process. By solving each simpler problem independently within a time step and subsequently recombining them, operator splitting method provides an approximate solution to the original problem (see reviews [35, 8]). This method is particularly advantageous for problems in which multiple processes, such as diffusion, advection, and reaction, interact simultaneously, as operator splitting method allows these processes to be separated and then treated with specialized methods optimized for each.
@@
{$&u^{\prime}(t)=A\left(u(t)\right)+B\left(u(t)\right),\quad t\in[0,T],\\ &u(0)~{}\text{given},\end{aligned}\right.$ (1.1)
@@
with the evolution operator $\mathcal{E}_{A+B}$. The evolutionary problem (1.1) is subdivided into two problems $u^{\prime}=A(u)$ and $u^{\prime}=B(u)$, whose solutions correspond to evolution operators $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$, respectively. By alternately employing these two evolution operators $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$, a $k$th-order splitting operator $\mathcal{S}^{[k]}$ can be systematically generated. Let $t_{n}=n\tau,~{}N\geq 0$ with the step size $\tau=T/N$. The corresponding splitting method is written as
$u^{n}=\mathcal{S}^{[k]}(\tau)u^{n-1}\approx u(t_{n})=\mathcal{E}_{A+B}(\tau)u(t_{n-1}),\quad 1\leq n\leq N$ (1.3)
that approximates the exact solution $u(t_{n})$ to (1.1). Typically, there are two forms of splitting operator $\mathcal{S}^{[k]}$: single-product expansion (SPE)
@@
and multi-product expansion (MPE)
@@
The earliest developments in operator splitting methods began with two SPE splittings: the first-order Lie–Trotter splitting
$\mathcal{S}^{[1],1}(\tau)=\mathcal{E}_{A}(\tau)\mathcal{E}_{B}(\tau)\quad\mbox{or}\quad\mathcal{S}^{[1],2}(\tau)=\mathcal{E}_{B}(\tau)\mathcal{E}_{A}(\tau),$ (1.6)
@@
$\mathcal{S}^{[2],1}(\tau)=\mathcal{E}_{A}\left(\frac{\tau}{2}\right)\mathcal{E}_{B}(\tau)\mathcal{E}_{A}\left(\frac{\tau}{2}\right)\quad\mbox{or}\quad\mathcal{S}^{[2],2}(\tau)=\mathcal{E}_{B}\left(\frac{\tau}{2}\right)\mathcal{E}_{A}(\tau)\mathcal{E}_{B}\left(\frac{\tau}{2}\right).$ (1.7)
@@
Historically, a consensus has emerged regarding the development of splitting techniques based on the Lie product formula [37]. However, the precise origins of this formula remain somewhat ambiguous [17, 21]. Trotter later generalized it to self-adjoint linear operators in [49], leading to the first-order Lie–Trotter splitting method. The second-order Strang splitting method is constructed by composing the Lie–Trotter splitting and its adjoint with half time steps. It was first introduced in [43] as an alternative approach to solving multidimensional problems with one-dimensional operators. As mentioned in [8], the Strang splitting has been a popular integrator, as second-order methods often achieve the right balance between accuracy and complexity. The Lie–Trotter and Strang splitting methods, while describing the systems they approximate well [21], are limited to first- or second-order accuracy in certain practical applications [30, 44, 31] where higher accuracy might be required.
@@
One technique to derive high-order (third-order and higher) SPE splitting methods (1.4) is the composition of lower-order splitting methods. For example, one can utilize the Strang splitting as a basic method and impose the symmetry of step coefficients on them to construct high-order SPE splittings. In [18], it was shown that even-order splitting methods can be recursively constructed through compositions of the Strang splitting method. In [50], a series of time-symmetric even-order splitting methods with exact time reversibility such as triple jump [18] are derived recursively from the Strang splitting method by leveraging the property that even-order terms in the Baker–Campbell–Hausdorff (BCH) formula can be canceled. Analogously, one can develop higher-order splitting methods by composing the Lie–Trotter splitting method with its adjoint while leveraging order conditions on step coefficients. These methods have been examined in [13], which provides an overview of several third-order operator splitting methods. Beyond composition, step coefficients can also be determined using various order conditions summarized by Blanes et al. in [8] to construct higher-order SPE methods. A key advantage of SPE splittings is that the resulting algorithms can preserve structure, such as symplecticity for Hamiltonian systems. Despite the ability of SPE splittings to achieve arbitrarily high order and preserve structure for certain models, they face a fundamental challenge beyond the second order. As noted by Chin and Geiser in [16], high-order SPE splittings cannot be applied to time-irreversible or semi-group problems due to the exponentially growing number of operators with unavoidable negative step coefficients [42, 46]. Even for time-reversible systems where backward time integration is not problematic, the exponential growth on the number of force evaluations renders high-order symplectic integrators difficult to be derived and expensive to use, as demonstrated in [16].
@@
Complex step coefficients can be admitted in SPE splittings [45, 3]. For parabolic equations, the existence of negative coefficients in high-order SPE splittings leads to instability, while high-order SPE splittings with complex step coefficients provide a possible way to overcome this order barrier [12, 25]. To maintain the stability of the heat propagator, complex step coefficients with positive real parts are required. In [25], the composition technique is applied to construct high-order SPE splitting methods with complex step coefficients. However, as shown in [6], if one considers complex step coefficients with positive real parts for SPE splittings, there are still order barriers for the composition technique based on the Lie–Trotter and Strang splittings. Meanwhile, in [4], a symmetric-conjugate method applied to linear unitary problems has been developed, which can achieve arbitrarily high order for several parabolic PDEs [12, 25] because these methods of order higher than two do exist with coefficients having positive real parts. However, these methods exhibit severe instabilities for the time-dependent Schrödinger equation, as both positive and negative imaginary parts will appear before the unbounded Laplace operators.
@@
The MPE splitting (1.5) generalizes the SPE splitting via linear combination, which enables the construction of arbitrarily high-order splitting methods without complex or negative step coefficients. A second-order MPE splitting, first mentioned by Strang in [43] and named symmetrically weighted sequential splitting in [19]:
@@
is derived from a linear combination of two Lie–Trotter splitting methods. In [15, 16], an efficient approach is presented for constructing high-order integrators based on time-symmetric splitting methods (for example, Strang splitting) and Richardson extrapolation. This approach not only preserves the geometric properties of the system (e.g., symplecticity for Hamiltonian problems), but also supports the construction of arbitrarily high-order splitting methods [9, 14, 5]. However, as rigorously shown by Suzuki and Sheng in [42, 46] and further strengthened in [22], regardless of how many operators are subdivided, any MPE splitting method of order higher than two inevitably requires at least one negative step coefficient or weight per operator. Since negative step coefficients lead to instability of some operators, we consider the MPE splitting method with all positive step coefficients, which implies the inevitable existence of negative weights.
@@
Despite numerous numerical experiments and practical applications that demonstrate the remarkable performance of high-order MPE splitting methods with all positive step coefficients [38, 16, 1], a rigorous analysis of their stability remains lacking. The primary challenge lies in the fact that trigonometric inequalities cannot be directly applied for estimation, as their upper bounds grow exponentially with iterations. In [39], the author suggested that the safest approach to designing high-order operator splitting methods involves negative weight coefficients while ensuring the positivity of step coefficients. In the book [26], Hundsdorfer and Verwer stated that “it is not precisely known for what kinds of problems MPE splitting methods with negative weights will be stable or unstable”. The stability analysis of MPE splitting methods with negative weights remains an open problem.
@@
In this work, we establish the uniform stability and consequently sharp convergence for arbitrarily high-order MPE splitting methods. For simplicity, we consider the semilinear parabolic equation with periodic boundary conditions as a typical problem. Our investigation addresses two main aspects. Firstly, to establish stability, we reformulate MPE splitting as several components, each formed by the composition of the three evolution operators $\mathcal{E}_{A}(\tau)$, $\mathcal{E}_{B}(\tau)$, and $\mathcal{E}_{B}(\tau)-\mathcal{I}$. By leveraging the stability properties of these evolution operators, we then derive uniform bounds for arbitrarily high-order MPE splitting methods. Secondly, we employ Lie derivative calculus and bounds for iterated Lie-commutators [27, 34, 48] to analyze consistency. By combining stability and consistency, we rigorously prove convergence by deriving a uniform Sobolev bound for the exact solution to the semilinear parabolic equation. In addition, we perform extensive numerical experiments to validate the accuracy, stability, and adaptive time-stepping strategies of these high-order MPE splitting methods, highlighting their efficiency and robustness through a range of nonlinear semilinear parabolic models.
@@
The paper is organized as follows. In Section 2, we introduce the basic definitions and the calculus of Lie derivatives utilized in the derivation of a local error expansion for generic MPE splitting methods. Section 3 summarizes existing formulations of MPE splittings and their construction methodologies, introduces a new perspective for their design, and proposes two new fourth-order splittings as illustrative examples based on the simplification techniques of order conditions. Section 4 states the semilinear parabolic equation as a typical equation and provides the first time a uniform stability result for generic MPE splitting methods. In Section 5, the consistency is established through the derivation of the local error expansion for arbitrarily high-order MPE splitting methods, followed by the proof of convergence. In Section 6, we conduct several numerical experiments that validate our theoretical results, demonstrating the stability and accuracy of the proposed high-order MPE splitting methods in various test cases. Finally, we conclude the paper in Section 7.
2 Notation and preliminaries
@@
Throughout this paper, we formally define the following notations. We denote by =$\mathbb{N}=\{m\in\mathbb{Z}:m\geq 0\}$ the set of non-negative integers. The compact multi-index notation $\mu=(\mu_{1},\ldots,\mu_{d})\in\mathbb{Z}^{d}$ and the vector notation $x=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$ are employed. Moreover, for $\alpha,\mu\in\mathbb{Z}^{d}$, relations such as $\leq$ are defined componentwise; that is, $\alpha\leq\mu$ iff $\alpha_{j}\leq\mu_{j}$ for all $1\leq j\leq d$. We set $|\mu|=\mu_{1}+\cdots+\mu_{d}$ and $\mu!=\mu_{1}!\cdots\mu_{d}!$ for $\mu\in\mathbb{N}^{d}$, and $c\mu=(c\mu_{1},\ldots,c\mu_{d})$ for $c\in\mathbb{Z}$ and $\mu\in\mathbb{Z}^{d}$. For $\mu\in\mathbb{N}^{d}$ and $x\in\mathbb{R}^{d}$, we write $x^{\mu}=x_{1}^{\mu_{1}}\cdots x_{d}^{\mu_{d}}$ and $\partial_{x}^{\mu}=\partial_{x_{1}}^{\mu_{1}}\cdots\partial_{x_{d}}^{\mu_{d}}$ for brevity and further denote by $\Delta=\partial_{x_{1}}^{2}+\cdots+\partial_{x_{d}}^{2}$ the $d$-dimensional Laplace operator. As standards, for $1\leq p\leq\infty$, we denote by $L^{p}(\Omega)$ the Lebesgue space of real-valued functions on a suitable domain $\Omega\subset\mathbb{R}^{d}$, which forms a Banach space with the associated norm given by
$\|u\|_{p}=\|u\|_{L^{p}(\Omega)}=\left({\int_{\Omega}|u(x)|^{p}\,\mathrm{d}x}\right)^{\frac{1}{p}},\quad\text{for}~{}~{}1\leq p<\infty,$ (2.1)
and
$\|u\|_{\infty}=\|u\|_{L^{\infty}(\Omega)}=\text{ess}\sup\{|u(x)|:x\in\Omega\},\quad\text{for}~{}~{}p=\infty.$ (2.2)
The Sobolev space $W^{m,p}(\Omega)$ consists of functions whose partial derivatives up to order $m\geq 1$ belong to $L^{p}(\Omega)$. In particular, the Hilbert spaces with periodic boundary conditions are $H^{0}_{\text{per}}(\Omega)=L^{2}(\Omega)$ and $H^{m}_{\text{per}}(\Omega)=W^{m,2}_{\text{per}}(\Omega)$ for $m\geq 1$, with the associated norm given by
@@
for brevity. Furthermore, for a family of (unbounded nonlinear) operators $\left(F_{\ell}\right)_{j\leq\ell\leq k}$ the product is defined downward (on a suitably chosen domain), that is, we set
$\prod_{\ell=j}^{k}F_{\ell}=F_{k}\cdots F_{j},\quad\text{if}\quad j\leq k,$ (2.4)
and $\prod_{\ell=j}^{k}F_{\ell}=\mathcal{I}$ if $j>k$ with $\mathcal{I}$ the identity operator. We make the convention that $|\Omega|^{\frac{1}{\infty}}=1$ for the bounded domain $\Omega$.
We employ the calculus of Lie derivatives (see, e.g., [24, Sect.III.5], [26, Sect.IX.1.4], or [28, Appendix.A]). Consider an initial value problem of the form
$u^{\prime}(t)=F(u(t)),\quad 0\leq t\leq T,\quad u(0)=u^{0},$ (2.5)
$u(t)=\mathcal{E}_{F}(t)u^{0},\quad 0\leq t\leq T,$ (2.6)
$\mathrm{e}^{tD_{F}}G(u^{0})=G(\mathcal{E}_{F}(t)u^{0}),\quad 0\leq t\leq T.$ (2.8)
$\mathrm{e}^{(t_{1}+t_{2})D_{F}}=\mathrm{e}^{t_{1}D_{F}}\mathrm{e}^{t_{2}D_{F}}=\mathrm{e}^{t_{2}D_{F}}\mathrm{e}^{t_{1}D_{F}},\quad 0\leq t_{1}+t_{2}\leq T,\left.\quad\mathrm{e}^{t_{1}D_{F}}\right|_{t_{1}=0}=\mathcal{I},$ (2.9)
$\mathrm{e}^{t_{1}D_{F_{1}}}\mathrm{e}^{t_{2}D_{F_{2}}}G(u^{0})=G\left(\mathcal{E}_{F_{2}}(t_{2})(\mathcal{E}_{F_{1}}(t)u^{0})\right),\quad 0\leq t_{1}+t_{2}\leq T.$ (2.10)
Additionally, we have the rule
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{e}^{tD_{F}}G(u^{0})=\mathrm{e}^{tD_{F}}D_{F}G(u^{0})=D_{F}\mathrm{e}^{tD_{F}}G(u^{0}),\quad 0\leq t\leq T.$ (2.11)
We also note that the identity
$\mathrm{e}^{tD_{F}}=\mathcal{I}+\left.\mathrm{e}^{\sigma D_{F}}\right|_{\sigma=0}^{t}=\mathcal{I}+\int_{0}^{t}\frac{\mathrm{~{}d}}{\mathrm{~{}d}\sigma}\mathrm{e}^{\sigma D_{F}}\mathrm{~{}d}\sigma=\mathcal{I}+\int_{0}^{t}\mathrm{e}^{\sigma D_{F}}D_{F}\mathrm{~{}d}\sigma,\quad 0\leq t\leq T,$ (2.12)
$\mathrm{e}^{tD_{F}}=\sum_{j=0}^{s-1}\frac{1}{j!}t^{j}D_{F}^{j}+\int_{T_{s}}\mathrm{e}^{\sigma_{s}D_{F}}D_{F}^{s}\mathrm{~{}d}\sigma,\quad 0\leq t\leq T,\quad s\in\mathbb{N}^{+},$ (2.13)
$\varphi_{0}(tD_{F})=\mathrm{e}^{tD_{F}},\quad\varphi_{j}(tD_{F})=\frac{1}{(j-1)!}\int_{0}^{1}\sigma^{j-1}\mathrm{e}^{(1-\sigma)tD_{F}}\mathrm{~{}d}\sigma,$ (2.14)
with the recurrence relation
$\varphi_{j}\left(tD_{F}\right)=\frac{1}{j!}\mathcal{I}+\varphi_{j+1}\left(tD_{F}\right)tD_{F},\quad j\geq 0.$ (2.15)
The Lie-commutator of two nonlinear operators $F,G$ is given by $\operatorname{ad}_{F}(G)(v)=[F,G](v)=F^{\prime}(v)G(v)-G^{\prime}(v)F(v)$; in particular, when $F$ and $G$ are linear, the above relation reduces to $\operatorname{ad}_{F}(G)=FG-GF$. In accordance with the above definition, we further set $\operatorname{ad}_{D_{F}}\left(D_{G}\right)(v)=\left[D_{F},D_{G}\right](v)=D_{F}D_{G}(v)-D_{G}D_{F}(v)=G^{\prime}(v)F(v)-F^{\prime}(v)G(v)$. By definition, the commutator $[D_{F},D_{G}]=-D_{[F,G]}.$ More generally, the iterated Lie-commutators are defined by induction
$\operatorname{ad}_{D_{F}}^{0}\left(D_{G}\right)=D_{G},\quad\operatorname{ad}_{D_{F}}^{j}\left(D_{G}\right)=\left[D_{F},\operatorname{ad}_{D_{F}}^{j-1}\left(D_{G}\right)\right],\quad j\geq 1,$ (2.16)
and
$\operatorname{ad}_{D_{F}}^{j}\left(D_{G}\right)=(-1)^{j}D_{\operatorname{ad}_{{F}}^{j}\left({G}\right)},\quad j\geq 1.$ (2.17)
3 MPE splitting methods for general evolutionary problems

### v2

@@ Abstract
The operator splitting method has been widely used to solve differential equations by splitting the equation into more manageable parts. In this work, we resolves a long-standing problem—how to establish the stability of multi-product expansion (MPE) splitting methods with negative weights. The difficulty occurs because negative weights in high-order MPE method cause the sum of the absolute values of weights larger than one, making standard stability proofs fail. In particular, we take the semilinear parabolic equation as a typical model and establish the stability of arbitrarily high-order MPE splitting methods with positive time steps but possibly negative weights. Rigorous convergence analysis is subsequently obtained from the stability result. Several numerical experiments validate the stability and accuracy of various high-order MPE splitting methods, highlighting their efficiency and robustness.
@@ 1 Introduction
Operator splitting method plays a critical role in approximating solutions of differential equations involving the coupling of multiple physical fields. The core idea is to subdivide the original equations into a series of simpler problems, each corresponding to a different physical or mathematical process. By solving each simpler problem independently within a time step and subsequently recombining them, operator splitting method provides an approximate solution to the original problem (see reviews [36, 8]). This method is particularly advantageous for problems in which multiple processes, such as diffusion, advection, and reaction, interact simultaneously, as operator splitting method allows these processes to be separated and then treated with specialized methods optimized for each.
@@
$\left\{\begin{aligned} &u^{\prime}(t)=A\left(u(t)\right)+B\left(u(t)\right),\qquad t\in[0,T],\\ &u(0)~\text{given},\end{aligned}\right$. (1.1)
@@
with the evolution operator $\mathcal{E}_{A+B}$. The evolutionary problem (1.1) is subdivided into two problems $u^{\prime}=A(u)$ and $u^{\prime}=B(u)$, whose solutions correspond to evolution operators $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$, respectively. By alternately employing these two evolution operators $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$, a $k$th-order splitting operator $\mathcal{S}^{[k]}$ can be systematically generated. Let $t_{n}=n\tau$ for $n=0,1,\cdots,N$, where the step size $\tau=T/N$ with $N\geq 1$. The corresponding splitting method is written as
$u^{n}=\mathcal{S}^{[k]}(\tau)u^{n-1}\approx u(t_{n})=\mathcal{E}_{A+B}(\tau)u(t_{n-1}),\qquad 1\leq n\leq N$ (1.3)
that approximates the exact solution $u(t_{n})$ to (1.1). Typically, the splitting operator $\mathcal{S}^{[k]}$ is composed by a single-product of the two evolution operators $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$:
@@
and multi-product expansion (MPE) splitting method is a linear combination of the splitting operators:
@@
The earliest developments in operator splitting methods began with two splittings: the first-order Lie–Trotter splitting
$\mathcal{S}^{[1],1}(\tau)=\mathcal{E}_{A}(\tau)\mathcal{E}_{B}(\tau)\qquad\mbox{or}\qquad\mathcal{S}^{[1],2}(\tau)=\mathcal{E}_{B}(\tau)\mathcal{E}_{A}(\tau),$ (1.6)
@@
$\mathcal{S}^{[2],1}(\tau)=\mathcal{E}_{A}\left(\frac{\tau}{2}\right)\mathcal{E}_{B}(\tau)\mathcal{E}_{A}\left(\frac{\tau}{2}\right)\qquad\mbox{or}\qquad\mathcal{S}^{[2],2}(\tau)=\mathcal{E}_{B}\left(\frac{\tau}{2}\right)\mathcal{E}_{A}(\tau)\mathcal{E}_{B}\left(\frac{\tau}{2}\right).$ (1.7)
@@
Historically, a consensus has emerged regarding the development of splitting techniques based on the Lie product formula [37]. However, the precise origins of this formula remain somewhat ambiguous [18, 24, 8, 36]. Trotter later generalized it to self-adjoint linear operators in [49], leading to the first-order Lie–Trotter splitting method. The second-order Strang splitting method is constructed by composing the Lie–Trotter splitting and its adjoint with half time steps. It was first introduced in [43] as an alternative approach to solving multidimensional problems with one-dimensional operators. As mentioned in [8], the Strang splitting has been a popular integrator, as second-order methods often achieve the right balance between accuracy and complexity. In some practical applications [33, 44, 32], higher accuracy might be required, whereas the Lie–Trotter and Strang splitting methods, while describing the systems well [24], are limited to first- and second-order accuracy.
@@
One technique to derive high-order (third-order and higher) splitting methods (1.4) is the composition of lower-order splitting methods. For example, one can utilize the Strang splitting as a basic method and impose the symmetry of step coefficients on them to construct high-order splitting methods. In [19], it was shown that even-order splitting methods can be recursively constructed through compositions of the Strang splitting method. In [50], a series of time-symmetric even-order splitting methods with exact time reversibility such as triple jump [19] are derived recursively from the Strang splitting method by leveraging the property that even-order terms in the Baker–Campbell–Hausdorff (BCH) formula can be canceled. Analogously, one can develop higher-order splitting methods by composing the Lie–Trotter splitting method with its adjoint while leveraging order conditions on step coefficients. These methods have been examined in [14], which provides an overview of several third-order operator splitting methods. Beyond composition, step coefficients can also be determined using various order conditions summarized by Blanes et al. in [8] to construct higher-order splitting methods. In addition, a broader range of high-order splitting methods is comprehensively summarized in [8, 36]. A key advantage of splitting methods is that the resulting algorithms can preserve structure, such as symplecticity for Hamiltonian systems. Despite the ability of splitting methods to achieve arbitrarily high order and preserve structure for certain models, they face a fundamental challenge beyond the second order. As noted in [13, 28], high-order splitting methods cannot be applied to time-irreversible or semigroup problems due to the exponentially growing number of operators with unavoidable negative step coefficients [42, 46]. Even for time-reversible systems where backward time integration is not problematic, the exponential growth on the number of force evaluations renders high-order symplectic integrators difficult to be derived and expensive to use, as demonstrated in [16].
@@
Complex step coefficients can be admitted in splitting methods [45, 2]. For parabolic equations, the existence of negative coefficients in high-order splitting methods leads to instability, while high-order splitting methods with complex step coefficients provide a possible way to overcome this order barrier [13, 28]. To maintain the stability of the heat propagator, complex step coefficients with positive real parts are required. In [28], the composition technique is applied to construct high-order splitting methods with complex step coefficients. However, as shown in [6], if one considers complex step coefficients with positive real parts for splitting methods, there are still order barriers for the composition technique based on the Lie–Trotter and Strang splittings. Meanwhile, in [3], a symmetric-conjugate method applied to linear unitary problems has been developed, which can achieve arbitrarily high order for several parabolic PDEs [13, 28] because these methods of order higher than two do exist with coefficients having positive real parts.
@@
The MPE splitting (1.5) generalizes the typical splitting methods via linear combination, which enables the construction of arbitrarily high-order splitting methods without complex or negative step coefficients. A second-order MPE splitting, first mentioned by Strang in [43] and named symmetrically weighted sequential splitting in [20]:
@@
is derived from a linear combination of two Lie–Trotter splitting methods. In [17, 16], an efficient approach is presented for constructing high-order integrators based on time-symmetric splitting methods (for example, Strang splitting) and Richardson extrapolation. Generally, MPE splitting methods do not preserve the geometric properties, since they are linear combinations of splitting methods. In certain cases, they can preserves the geometric properties (e.g., symplecticity for Hamiltonian problems) and achieve arbitrarily high order [5, 15, 10]. However, as rigorously shown by Suzuki and Sheng in [42, 46] and further strengthened in [25], regardless of how many operators are subdivided, any MPE splitting method of order higher than two inevitably requires at least one negative step coefficient or weight per operator. Since negative step coefficients lead to instability of some operators, we consider the MPE splitting method with all positive step coefficients, which implies the inevitable existence of negative weights. In other words, the MPE splitting methods we are concerned with are linear combinations of splitting methods up to order 2, with at least one negative weight.
@@
Despite numerous numerical experiments and practical applications that demonstrate the remarkable performance of high-order MPE splitting methods with all positive step coefficients [38, 16, 1], a rigorous analysis of their stability remains lacking. The primary challenge lies in the fact that trigonometric inequalities cannot be directly applied for estimation, as their upper bounds grow exponentially with iterations. In [39], the author suggested that the safest approach to designing high-order operator splitting methods involves negative weight coefficients while ensuring the positivity of step coefficients. In the book [29, Section 1.3], Hundsdorfer and Verwer stated that “it is not precisely known for what kinds of problems this scheme (a MPE splitting with negative weights) will be stable or unstable”. The stability analysis of MPE splitting methods with negative weights remains an open problem.
@@
In this work, we establish the stability and consequently sharp convergence for arbitrarily high-order MPE splitting methods. For simplicity, we consider the semilinear parabolic equation with periodic boundary conditions as a typical problem. Our investigation addresses two main aspects. Firstly, to establish stability, we reformulate MPE splitting as several components, each formed by the composition of the three evolution operators $\mathcal{E}_{A}(\tau)$, $\mathcal{E}_{B}(\tau)$, and $\mathcal{E}_{B}(\tau)$-$\mathcal{I}$. By leveraging the stability properties of these evolution operators, we then derive Sobolev bounds for arbitrarily high-order MPE splitting methods. Secondly, we employ Lie derivative calculus and bounds for iterated Lie-commutators [30, 35, 48] to analyze consistency. By combining stability and consistency, we rigorously prove convergence by deriving a Sobolev bound for the exact solution to the semilinear parabolic equation. In addition, we perform several numerical experiments to validate the accuracy, stability, and adaptive time-stepping strategies of these high-order MPE splitting methods, highlighting their efficiency and robustness through a range of nonlinear semilinear parabolic models.
@@
The paper is organized as follows. In Section 2, we introduce some notations and preliminaries. Section 3 summarizes existing formulations of MPE splittings and proposes two new fourth-order splittings. Section 4 states the semilinear parabolic equation as a typical equation and provides for the first time a stability result for generic MPE splitting methods. In Section 5, the consistency is established through the derivation of the local error expansion for arbitrarily high-order MPE splitting methods, followed by the proof of convergence. In Section 6, we conduct several numerical experiments that validate our theoretical results, demonstrating the stability and accuracy of the proposed high-order MPE splitting methods in various test cases. Finally, we conclude the paper in Section 7.
2 Notations and preliminaries
@@
Throughout this paper, we formally define the following notations. We denote by $\mathbb{N}=\{m\in\mathbb{Z}:m\geq 0\}$ the set of non-negative integers. The compact multi-index notation $\mu=(\mu_{1},\ldots,\mu_{d})\in\mathbb{Z}^{d}$ and the vector notation $x=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$ are employed. Moreover, for $\alpha,\mu\in\mathbb{Z}^{d}$, relations such as $\leq$ are defined componentwise; that is, $\alpha\leq\mu$ iff $\alpha_{j}\leq\mu_{j}$ for all $1\leq j\leq d$. We set $|\mu|=\mu_{1}+\cdots+\mu_{d}$ and $\mu!=\mu_{1}!\cdots\mu_{d}!$ for $\mu\in\mathbb{N}^{d}$, and $c\mu=(c\mu_{1},\ldots,c\mu_{d})$ for $c\in\mathbb{Z}$ and $\mu\in\mathbb{Z}^{d}$. For $\mu\in\mathbb{N}^{d}$ and $x\in\mathbb{R}^{d}$, we write $x^{\mu}=x_{1}^{\mu_{1}}\cdots x_{d}^{\mu_{d}}$ and $\partial_{x}^{\mu}=\partial_{x_{1}}^{\mu_{1}}\cdots\partial_{x_{d}}^{\mu_{d}}$ for brevity and further denote by $\Delta=\partial_{x_{1}}^{2}+\cdots+\partial_{x_{d}}^{2}$ the $d$-dimensional Laplace operator. The Sobolev space $W^{m,p}(\Omega)$ consists of functions whose partial derivatives up to order $m\geq 1$ belong to $L^{p}(\Omega)$. In particular, the Hilbert spaces with periodic boundary conditions are $H^{0}_{\text{per}}(\Omega)=L^{2}(\Omega)$ and $H^{m}_{\text{per}}(\Omega)=W^{m,2}_{\text{per}}(\Omega)$ for $m\geq 1$, with the associated norm given by
@@
for brevity. We make the convention that $|\Omega|^{\frac{1}{\infty}}=1$ where $|\Omega|$ is the volume of the bounded domain $\Omega$.
Furthermore, for a family of (unbounded nonlinear) operators $\left(F_{\ell}\right)_{j\leq\ell\leq k}$ the product is defined downward (on a suitably chosen domain), that is, we set
$\prod_{\ell=j}^{k}F_{\ell}=F_{k}\cdots F_{j},\qquad\text{if}\qquad j\leq k,$ (2.2)
and $\prod_{\ell=j}^{k}F_{\ell}=\mathcal{I}$ if $j>k$ with $\mathcal{I}$ the identity operator.
3 Some MPE splitting methods